The population mean and standard deviation are given below. Find the required probability and determine whether the given sample mean would be considered unusual.

For a sample of

nequals=60

find the probability of a sample mean being less than

24.1

if

muμequals=24

sigma=1.33
Answer #1

Sol:

xbar follows normal distibution with mean=xbar=24

s=sigma/sqrt(n)=1.33/sqrt(60)=0.1717023

P(Xbar<24.1)

z=xbar-mu/s

=24.1-24/0.1717023

=0.1/0.1717023

z= 0.5824034

P(Z< 0.5824034)

=0.7199

ANSWER:

the probability of a sample mean being less than 24.1=0.7199

it is not unusual